# Hierarchical Correlation Clustering and Tree Preserving Embedding

### v2

@@ 1 Introduction
Data clustering plays an essential role in unsupervised learning and exploratory data analysis. It is used in a variety of applications including web mining, network analysis, user understanding, image segmentation, knowledge management, and bioinformatics. Its goal is to partition the data into groups in a way that the objects in the same cluster are more similar according to some criterion, compared to the objects in different clusters.
@@
Some clustering methods partition the data into $K$ flat clusters for example via minimizing a cost function, e.g., $K$-means [51], spectral clustering [64, 58] and correlation clustering [9]. In many applications, however, the clusters are desired to be presented at different levels, such that both high-level and detailed information are preserved. Therefore, hierarchical clustering is useful to produce such structures, usually encoded by a dendrogram. A dendrogram is a tree data structure wherein every node represents a cluster and the leaf nodes (i.e., the nodes linked to just one other node) include only one object. A cluster at a higher level consists of the union of the lower-level clusters, and the inter-cluster dissimilarity between them.
@@
Hierarchical clustering can be performed either in an agglomerative (i.e., bottom-up) or in a divisive (i.e., top-down) manner [52]. Agglomerative methods are often computationally more efficient, making them more popular in practice [60]. In both approaches, the clusters are combined or divided according to different criteria, e.g., single, complete, average, centroid and Ward. Several methods aim to improve these methods. The work in [3] studies the locality and outer consistency of agglomerative algorithms in an axiomatic way. The works in [45, 48] focus on the statistical significance of hierarchical clustering. [21, 22, 61] investigate optimization methods for hierarchical clustering and propose approximate solutions. K-Linkage in [76] takes into account multiple dissimilarities for a pair of clusters. [7] employs global information to determine the inter-cluster similarities in order to eliminate the influence of noisy similarities, and [16] suggests using agglomerative methods on small dense subsets of the data, instead of the original data objects. The methods in [29, 34] investigate combining agglomerative methods with probabilistic models which then yields an extra computational complexity. Finally, [56, 20] develop efficient but approximate implementations of agglomerative methods.
@@
On the other hand, most clustering methods, either flat or hierarchical, assume non-negative pairwise (dis)similarities. However, in several practical applications, pairwise relations are provided in the form of both positive and negative similarities. For example, it might be more convenient for a user/oracle to specify if two objects are similar (i.e., a positive relation) or dissimilar (i.e., a negative relation) instead of giving only a positive (non-negative) pairwise similarity even if the two objects are dissimilar. In addition, the former provides more accurate information, as in the latter case, the dissimilarity between two objects (i.e., zero similarity) might be confused with the lack of information. Assuming the pairwise similarities are provided by an oracle, then uncertainty can be handled better by this kind of data, where the oracle can explicitly distinguish between noise, uncertainty, and lack of similarity. Some relevant applications for this setting include image segmentation with higher order correlation information [43, 44], community detection over graphs [63], web people search via analysis of connections [39], working with attraction/rejection data [23], automated label generation from clicks [4], webpage segmentation [12], social media mining [69], and entity resolution [8, 30].
@@
Thereby, a specific form of clustering model, called correlation clustering, has been developed to deal with such data. This model was first introduced on the graphs with only $+1$ or $-1$ pairwise similarities [9, 8], and then was generalized to the graphs with arbitrary positive or negative edge weights [23, 6, 13]. The original model obtains the number of clusters automatically. The variant in [17] limits the number of clusters to fixed $K$ clusters. Semidefinite programming (SDP) relaxation provides tight approximation bounds in particular for maximizing the agreements [13, 54], although it is computationally inefficient in practice [70]. Then, the methods in [15, 70] provide efficient greedy algorithms based on local search and Frank-Wolfe optimization with a fast convergence rate.
@@
However, all of these methods produce flat correlation clusters. In this paper, we first propose a Hierarchical Correlation Clustering (HCC) method that handles both positive and negative pairwise (dis)similarities and produces clusters at different levels (Section 3). To the best of our knowledge, this work is the first generalization of the well-studied correlation clustering to hierarchical clustering. [^1] We note the so-called hierarchical correlation clustering methods proposed in [2, 32, 49] are irrelevant to the well-studied correlation clustering problem [9, 8]; they study for example the correlation coefficients for high-dimensional data.
@@
One way to perform representation learning from hierarchical clustering is to compute an embedding that corresponds to the respective hierarchy. Tree preserving embedding [65, 66] is a method that achieves this for the special case of single linkage method. Later, [18] develops tree preserving embedding for various standard agglomerative clustering methods. We then adapt these works to develop a tree preserving embedding for HCC dendrograms (Section 4) where the embedded features can be used as a set of new features for an arbitrary downstream task. In this way, we can investigate HCC for the purpose of computing relevant features for a probabilistic method such as Gaussian Mixture Model (GMM), instead of solely using HCC for the purpose of hierarchical clustering. This enables us to apply a method like GMM for clustering the pairwise similarities that can be positive or negative numbers, a task that was not possible before.
Another representation learning paradigm that we study is called minimax dissimilarity, a graph-based method that is tightly connected to hierarchical clustering. It provides a sophisticated way to infer transitive relations and extract manifolds and elongated clusters in an unsupervised way [42, 14, 50]. Thereby, for the first time, we study minimax dissimilarities on the graphs with positive and negative (dis)similarities, i.e., with correlation clustering (Section 5). We show that using minimax dissimilarities with correlation clustering not only helps for extracting elongated patterns, but also yields a significant reduction in the computational complexity, i.e., from NP-hardness to a polynomial runtime.
@@ 2 Notations and Definitions
A dataset is specified by a set of $n$ objects with indices $\mathbf{O}=\{1,...,n\}$ and a pairwise similarity or dissimilarity matrix. An $n\times n$ matrix $\mathbf{S}\in\mathbb{R}^{n\times n}$ represents the pairwise similarities between the objects, whereas, the pairwise dissimilarities are shown by matrix $\mathbf{D}\in\mathbb{R}^{n\times n}$. Both of similarities and dissimilarities can be positive or negative. This property allows us to convert the pairwise similarities to dissimilarities by a simple transformation such as $\mathbf{D}=-\mathbf{S}$.  [^2] Such a nonparametric transformation resolves the issues related to obtaining a proper similarity measure from pairwise dissimilarities. For example, with kernels, e.g., RBF kernels, finding the optimal parameter(s) is often crucial and nontrivial, and the optimal parameters occur inside a very narrow range [57, 74]. Moreover, the methods we develop in this paper are unaffected by the choice of the transformation $\mathbf{D}$; for example in Algorithm 1, we only use the pairwise similarities $\mathbf{S}$. The objects and the pairwise (dis)similarities are represented by graph $\mathcal{G}=(\mathbf{O},\mathbf{S})$ or $\mathcal{G}=(\mathbf{O},\mathbf{D})$.
@@
The clusters at the lowest level, called leaf clusters/node, are the individual distinct objects, i.e., $\mathbf{v}$ is a leaf cluster if and only if $|\mathbf{v}|=1$. A cluster at a higher level contains the union of the objects of its children. The root of a dendrogram is defined as the cluster at the highest level which has the maximum size, i.e., all other clusters are its descendants. $linkage(\mathbf{v}),\mathbf{v}\in T$ returns the dissimilarity between the children of $\mathbf{v}$ based on the criterion used to compute the dendrogram (i.e., $dis(c_{l},c_{r})$ where $c_{l}$ and $c_{r}$ indicate the two child clusters of $\mathbf{v}$). For simplicity of explanation, w.l.g., we sometimes assume every non-leaf cluster has two children. The level of cluster $\mathbf{v}$, i.e., $level(\mathbf{v})$, is determined by
@@ 3 Hierarchical Correlation Clustering
Single linkage, complete linkage and average linkage methods are invariant w.r.t. the shift of the pairwise dissimilarities by constant $\alpha$.
Thus, we can still use these methods even with possibly negative pairwise dissimilarities as shifting the pairwise dissimilarities (by a sufficiently large constant) to make them non-negative does not change the solution. However, clustering the data consisting of positive and negative dissimilarities is usually conducted by correlation clustering. Thus, despite the applicability of single linkage, average linkage and complete linkage methods, we propose a novel hierarchical clustering consistent with the standard correlation clustering, called Hierarchical Correlation Clustering (HCC). This method is thus adapted to positive/negative pairwise (dis)similarities, and as our experiments confirm, it outperforms the other methods (i.e., SL, CL, and AL) when applied to such data.
@@
The cost function can be rewritten as
@@
$\;+\underbrace{\frac{1}{2}\sum_{k=1}^{K}\sum_{i,j\in\mathbf{v}_{k}}|\mathbf{S}_{ij}|+\frac{1}{2}\sum_{k=1}^{K}\sum_{\begin{subarray}{c}k^{\prime}=1,\\ k^{\prime}\neq k\end{subarray}}^{K}\sum_{i\in\mathbf{v}_{k}}\sum_{j\in\mathbf{v}_{k^{\prime}}}|\mathbf{S}_{ij}|}_{constant}$ (7)
@@
$R^{CC}(\mathbf{v}_{1},...,\mathbf{v}_{K};\mathbf{S})=constant+\sum_{k=1}^{K}\sum_{\begin{subarray}{c}k^{\prime}=1,\\ k^{\prime}\neq k\end{subarray}}^{K}\sum_{i\in\mathbf{v}_{k}}\sum_{j\in\mathbf{v}_{k^{\prime}}}\mathbf{S}_{ij}$ (15)
@@
$dis^{CC}(\mathbf{u},\mathbf{v})=\sum_{i\in\mathbf{u}}\sum_{j\in\mathbf{v}}\mathbf{D}_{ij}=-\sum_{i\in\mathbf{u}}\sum_{j\in\mathbf{v}}\mathbf{S}_{ij}$ (19)

### v4

@@ 1 Introduction
Data clustering plays an essential role in unsupervised learning and exploratory data analysis. It is used in a variety of applications including web mining, network analysis, image segmentation, bioinformatics, user analytics and knowledge management. Its goal is to partition the data into groups in a way that the objects in the same cluster are more similar according to some criterion, compared to the objects in different clusters.
@@
Many clustering methods partition the data into $K$ flat clusters for example, $K$-means [55], spectral clustering [68, 62] and correlation clustering [8]. In many applications, however, the clusters are preferred to be presented at different levels, encompassing both high-level and detailed information. Hierarchical clustering is useful to produce such structures, usually encoded by a dendrogram. A dendrogram is a tree data structure where each node corresponds to a cluster, with the leaf nodes (those at the bottom of the tree) containing only one object. Higher-level clusters are formed by aggregating lower-level clusters and the inter-cluster dissimilarity between them.
@@
Hierarchical clustering can be performed either in an agglomerative (i.e., bottom-up) or in a divisive (i.e., top-down) manner [56]. Agglomerative methods are often computationally more efficient, making them more popular in practice [64]. In both approaches, the clusters are aggregated or split based on various criteria, such as single, average, centroid, complete and Ward. Several studies aim to improve these methods. The works in [49, 52] focus on the statistical significance of hierarchical clustering. [24, 25, 65] formulate this problem as an optimization problem and propose approximate solutions. [82] considers multiple dissimilarities for a pair of clusters, and [11, 17] suggest merging multiple clusters at each step instead of one. [6] employs global information to eliminate the influence of noisy similarities, and [19] proposes to apply agglomerative methods to small subsets of the data instead of individual data objects. [33, 38] augment agglomerative methods with probabilistic models, and finally, [60, 23] propose efficient but approximate methods for hierarchical clustering.
@@
On the other hand, most clustering methods, either flat or hierarchical, assume non-negative pairwise (dis)similarities. However, in several practical applications, pairwise similarities can be any real number, positive or negative. For example, it could be preferable for a user or oracle to indicate whether two objects are similar (considered a positive relation) or dissimilar (considered a negative relation), rather than solely providing a positive (non-negative) pairwise similarity, even if the two objects are dissimilar. The former approach yields more precise information because, in the latter scenario, the dissimilarity between two objects (i.e., zero similarity) could be confused with a lack of available information. Some relevant applications for this setting include image segmentation with higher order correlation information [47, 48], webpage segmentation [12], community detection over graphs [67], social media mining [73], analysis of connections over web [43], dealing with attraction/rejection data [26], automated label generation from clicks [3] and entity resolution [7, 34].
@@
Hence, a specialized clustering model known as correlation clustering has been developed to work with such data. This model was first introduced on the graphs with only $+1$ or $-1$ pairwise similarities [8, 7], and then was generalized to the graphs with arbitrary positive or negative edge weights [26, 5, 13]. The original model obtains the number of clusters automatically. The variant in [20] limits the number of clusters to fixed $K$ clusters. Semidefinite programming (SDP) relaxation provides tight approximation bounds in particular for maximizing the agreements [13, 58], although it is computationally inefficient in practice [74]. Then, [15, 74] provide efficient greedy algorithms based on local search and Frank-Wolfe optimization with a fast convergence rate.
@@
However, all of these methods produce flat correlation clusters. In this paper, we first propose a Hierarchical Correlation Clustering (HCC) method that handles both positive and negative pairwise (dis)similarities and produces clusters at different levels (Section 3). To the best of our knowledge, this work is one the first extensions of the well-known correlation clustering to hierarchical clustering. [^1] We note the so-called hierarchical correlation clustering methods proposed in [2, 36, 53] are irrelevant to the well-studied correlation clustering problem [8, 7]; they study for example the correlation coefficients for high-dimensional data. A hierarchical correlation clustering, also called HCC, is developed in [76]. This method offers a $0.4767$ approximation, but lacking experimental evaluation. Furthermore, unlike our method, the method in [76] does not follow the generic agglomerative clustering procedure.
@@
One way to perform representation learning from hierarchical clustering is to compute an embedding that corresponds to the respective hierarchy. Tree preserving embedding [69, 70] is a method that achieves this for the special case of single linkage method. Later, [21] develops tree preserving embedding for various standard agglomerative clustering methods. We then adapt these works (in particular the later work [21]) to develop a tree preserving embedding for HCC dendrograms (Section 4) where the embedded features can be used as a set of new features for an arbitrary downstream task. In this way, we can investigate HCC for the purpose of computing relevant features for a probabilistic method such as Gaussian Mixture Model (GMM), instead of solely using HCC for the purpose of hierarchical clustering. This enables us to apply a method like GMM for clustering the pairwise similarities that can be positive or negative numbers, a task that was not possible before.
Another representation learning paradigm that we study is called minimax dissimilarity, a graph-based method that is tightly connected to hierarchical clustering. It provides a sophisticated way to infer transitive relations and extract manifolds and elongated clusters in an unsupervised way [32, 46, 14, 54]. Thereby, for the first time, we study minimax dissimilarities on the graphs with positive and negative (dis)similarities, i.e., with correlation clustering (Section 5). We show that using minimax dissimilarities with correlation clustering not only helps for extracting elongated patterns, but also yields a significant reduction in the computational complexity, i.e., from NP-hardness to a polynomial runtime.
@@ 2 Notations and Definitions
A dataset is characterized by a set of $n$ objects with indices $\mathbf{O}=\{1,...,n\}$ and a pairwise similarity or dissimilarity matrix. An $n\times n$ matrix $\mathbf{S}\in\mathbb{R}^{n\times n}$ represents the pairwise similarities between the objects, whereas, the pairwise dissimilarities are shown by matrix $\mathbf{D}\in\mathbb{R}^{n\times n}$. Both of similarities and dissimilarities can be positive or negative. This property allows us to convert the pairwise similarities to dissimilarities by a simple transformation such as $\mathbf{D}=-\mathbf{S}$, i.e., the pairwise dissimilarities are obtained by the negation of the similarities and vice versa.  [^2] Such a nonparametric transformation resolves the issues related to obtaining a proper similarity measure from pairwise dissimilarities. For example, with kernels, e.g., RBF kernels, finding the optimal parameter(s) is often crucial and nontrivial, and the optimal parameters occur inside a very narrow range [61, 80]. Moreover, the methods we develop in this paper are unaffected by the choice of the transformation $\mathbf{D}$; for example in Algorithm 1, we only use the pairwise similarities $\mathbf{S}$. The objects and the pairwise (dis)similarities are represented by graph $\mathcal{G}=(\mathbf{O},\mathbf{S})$ or $\mathcal{G}=(\mathbf{O},\mathbf{D})$.
@@
The clusters at the lowest level, called leaf clusters/node, are the individual distinct objects, i.e., $\mathbf{v}$ is a leaf cluster if and only if $|\mathbf{v}|=1$. A cluster at a higher level contains the union of the objects of its children. The root of a dendrogram is defined as the cluster at the highest level which has the maximum size, i.e., all other clusters are its descendants. $linkage(\mathbf{v}),\mathbf{v}\in T$ returns the dissimilarity between the children of $\mathbf{v}$ based on the criterion used to compute the dendrogram (i.e., $dis(c_{l},c_{r})$ where $c_{l}$ and $c_{r}$ indicate the two child clusters of $\mathbf{v}$). For simplicity of explanation, w.l.g., we assume every non-leaf cluster has two children. The level of cluster $\mathbf{v}$, i.e., $level(\mathbf{v})$, is determined by
@@ 3 Hierarchical Correlation Clustering
Single linkage, complete linkage and average linkage methods are invariant w.r.t. the shift of the pairwise dissimilarities by an arbitrary real number $\alpha$.
Thus, we can still use these methods even with possibly negative pairwise dissimilarities as shifting the pairwise dissimilarities (by a large enough constant) to make them non-negative does not change the solution.
However, clustering the data consisting of positive and negative dissimilarities is usually conducted by correlation clustering. Thus, despite the applicability of single linkage, average linkage and complete linkage methods, we propose a novel hierarchical clustering consistent with the standard correlation clustering, called Hierarchical Correlation Clustering (HCC). This method is thus adapted to positive/negative pairwise (dis)similarities, and as our experiments confirm, it outperforms the other methods (i.e., SL, CL, and AL) when applied to such data.
@@
We may rewrite the cost function as
@@
$\;+\underbrace{\frac{1}{2}\sum_{k=1}^{K}\sum_{i,j\in\mathbf{v}_{k}}|\mathbf{S}_{ij}|+\frac{1}{2}\sum_{k=1}^{K}\sum_{\begin{subarray}{c}k^{\prime}=1,\\ k^{\prime}\neq k\end{subarray}}^{K}\sum_{i\in\mathbf{v}_{k}}\sum_{j\in\mathbf{v}_{k^{\prime}}}|\mathbf{S}_{ij}|}_{constant}$
@@
$R^{CC}(\mathbf{v}_{1},...,\mathbf{v}_{K};\mathbf{S})=constant+\sum_{k=1}^{K}\sum_{\begin{subarray}{c}k^{\prime}=1,\\ k^{\prime}\neq k\end{subarray}}^{K}\sum_{i\in\mathbf{v}_{k}}\sum_{j\in\mathbf{v}_{k^{\prime}}}\mathbf{S}_{ij}$
@@
$dis^{CC}(\mathbf{u},\mathbf{v})=\sum_{i\in\mathbf{u}}\sum_{j\in\mathbf{v}}\mathbf{D}_{ij}=-\sum_{i\in\mathbf{u}}\sum_{j\in\mathbf{v}}\mathbf{S}_{ij}.$ (5)